# Non-local Hamilton–Jacobi-Bellman equations for the stochastic optimal control of path-dependent piecewise deterministic processes

### v1

@@ 5.1. The operator $G$
(i) Step 1 (construction of $\bar{\psi}$). Consider a partition $0=r_{0}<r_{1}<\ldots<r_{N}=T$ of $[0,T]$ with
$r_{k}-r_{k-1}<\frac{1}{2}\quad\textup{for all}\,\,k.$
Then $G_{r_{N-1},r_{N};h}$ is a contraction, say there is a $\kappa\in(0,1)$ such that, for all $\psi_{1}$, $\psi_{2}\in\mathrm{BLSA}([r_{N_{1}},r_{N}]\times\Omega)$, we have141414 If $\alpha\in\mathcal{A}$, then, for some probability measure ${{\mathbb{P}}}$, we have $[G_{r_{N-1},r_{N};h,\alpha}\,\psi_{1}](s,x)-[G_{r_{N-1},r_{N};h,\alpha}\,\psi_{2}](s,x)\leq\int_{r_{N-1}}^{r_{N}}{{\mathbb{E}}}^{{\mathbb{P}}}\lVert{\psi_{1}-\psi_{2}}\rVert_{\infty}\,{\mathrm{d}}t\leq\frac{1}{2}\lVert{\psi_{1}-\psi_{2}}\rVert_{\infty}.$ If $(s,x)\in[r_{N-1},r_{N}]\times\Omega$ and ${\varepsilon}>0$, then there is an $\alpha^{\varepsilon}\in\mathcal{A}$ with $[G_{r_{N-1}.r_{N};h}\,\psi_{1}](s,x)-[G_{r_{N-1}.r_{N};h}\,\psi_{2}](s,x)$ $\qquad\leq[G_{r_{N-1}.r_{N};h}\,\psi_{1}](s,x)-[G_{r_{N-1}.r_{N};h,\alpha^{\varepsilon}}\,\psi_{2}](s,x)+{\varepsilon}$ $\qquad\leq[G_{r_{N-1}.r_{N};h,\alpha^{\varepsilon}}\,\psi_{1}](s,x)-[G_{r_{N-1}.r_{N};h,\alpha^{\varepsilon}}\,\psi_{2}](s,x)+{\varepsilon}\leq\frac{1}{2}\lVert{\psi_{1}-\psi_{2}}\rVert_{\infty}+{\varepsilon}.$
$\lVert{G_{r_{N-1},r_{N};h}\psi_{1}-G_{r_{N-1},r_{N};h}\psi_{2}}\rVert_{\infty}\leq\kappa\lVert{\psi_{1}-\psi_{2}}\rVert_{\infty},$
and thus has a unique fixed point $\bar{\psi}_{N}$. Together with (4.6), we obtain [^15] Note at this point, we have not established yet that $V^{\prime}$ is lower semi-analytic.
$\leq\kappa\lVert{V^{\prime}|_{[r_{N-1},r_{N}]\times\Omega}-\bar{\psi}_{N}}\rVert_{\infty},$
i.e.,
$\bar{\psi}_{N}=V^{\prime}|_{[r_{N-1},r_{N}]\times\Omega}$
and thus $V^{\prime}$ is lower semi-analytic on $[r_{N-1},r_{N}]\times\Omega$.
Similarly, $G_{r_{N-2},r_{N-1};\bar{\psi}_{N}}$ is a contraction and thus has a unique fixed point $\bar{\psi}_{N-1}$, which coincides with $V^{\prime}$ on $[r_{N-2},r_{N-1}]\times\Omega$ thanks to Corollary 4.12.
Iterating this process yields a finite sequence $(\bar{\psi})_{k=1}^{N}$ with the properties that each $\bar{\psi}_{k}$ is the unique fixed point of $G_{r_{k-1},r_{k};\bar{\psi}_{k+1}}$, where we set
$\bar{\psi}_{N+1}(s,x):=h(x),$
and that $\bar{\psi}_{k}$ coincides with $V^{\prime}$ on $[r_{k-1},r_{k}]\times\Omega$ if $k>2$ and on $(0,r_{1}]\times\Omega$ if $k=1$.
Next, define $\bar{\psi}:[0,T]\times\Omega\to{\mathbb{R}}$ by
Step 2 ($\bar{\psi}$ is a fixed point of $G$). Consider the function $\bar{v}:[0,T]\times\Omega\to{\mathbb{R}}$ defined by
$\bar{v}(s,x):=[G\bar{\psi}](s,x).$
Then $\bar{v}=\bar{\psi}_{N}$ on $[r_{N-1},r_{N}]\times\Omega$.
$\bar{v}(s,x)=[(G_{r_{N-2},r_{N-1};\bar{v}})\bar{\psi}_{N-1}](s,x)=\bar{\psi}_{N-1}(s,x).$
Iterating this procedure yields $\bar{v}=\bar{\psi}$, i.e., $\bar{\psi}$ is a fixed point of $G$161616 Note this argument does not depend on properties of $V^{\prime}$ such as (4.6), which to establish was relatively technical. A simpler argument would be to just note that $V$ is lower semi-analytic by Step 1 and that then, together with (4.6), we can write $V=GV$.
Step 3 ($\bar{\psi}$ is the unique fixed point of $G$). Let $\bar{\psi}^{\prime}$ be a fixed point of $G$. On $[r_{N-1},r_{N}]\times\Omega$, we have
$\bar{\psi}^{\prime}=G\bar{\psi}^{\prime}=(G_{r_{N-1},r_{N};h})\bar{\psi}^{\prime}.$
Thus $\bar{\psi}^{\prime}$ coincides on $[r_{N-1},r_{N}]\times\Omega$ with $\bar{\psi}_{N}$, the unique fixed point of $G_{r_{N-1},r_{N};h}$.
Hence, $\bar{\psi}^{\prime}$ coincides on $[r_{N-2},r_{N-1}]\times\Omega$ with $\bar{\psi}_{N-1}$, the unique fixed point of $G_{r_{N-2},r_{N-1};\bar{\psi}_{N}}$. Iterating this process yields $\bar{\psi}^{\prime}=\bar{\psi}$.
Remark 5.5.
The fixed point $\bar{\psi}$ satisfies the dynamic programming principle, i.e., if $0\leq s\leq s_{1}\leq T$ and $x\in\Omega$, then
$\bar{\psi}(s,x)=\inf_{a\in\mathcal{A}}\Bigl{\{}\int_{s}^{s_{1}}\chi^{s,x,a}(t)\,\Bigl{[}\ell^{s,x,a}(t)$
$+\int_{{\mathbb{R}}^{d}}\bar{\psi}(t,\phi^{s,x,a}\otimes_{t}e)\,\lambda^{s,x,a}(t)\,Q^{s,x,a}(t,{\mathrm{d}}e)\Bigr{]}\,{\mathrm{d}}t+\chi^{s,x,a}(s_{1})\bar{\psi}(s_{1},\phi^{s,x,a})\Bigr{\}}.$
Next, we show that the fixed point $\bar{\psi}$ is sufficiently regular. To simplify notation, define $\lVert{\cdot}\rVert_{s}$, $s\in[0,T]$, on $\Omega$ by
$\lVert{x}\rVert_{s}:=\sup_{0\leq t\leq s}\left|{x(t)}\right|.$
Also put
Lemma 5.6.
for all $s\in[s_{1},s_{2}]$, $x$, $\tilde{x}\in\Omega$. Let $\lVert{\eta}\rVert_{\infty}\leq\lVert{\bar{\psi}}\rVert_{\infty}$. Recall $G_{s_{1},s_{2};\eta}$ from (5.4).
Then
for all $(s,x,\tilde{x})\in[s_{1},s_{2}]\times\Omega\times\Omega$.
Proof.
See section B.3. ∎
Lemma 5.7.
There is a constant $L_{\bar{\psi}}\in{\mathbb{R}}_{+}$ such that, for every $s\in[0,T]$, $x$, $\tilde{x}\in\Omega$,
Proof.
See section B.4. ∎
5.2. Proof of Theorem 5.2
Our approach can be considered as a path-dependent counterpart of [DavisFarid]. First, given $\psi\in\mathrm{BLSA}([0,T]\times\Omega)$, consider the terminal-value problem
($\text{{TVP}}_{\psi}$) $\begin{split}-\partial_{t}u(t,x)-F_{\psi}(t,x,u(t,x),\partial_{x}u(t,x))&=0,\quad(t,x)\in[0,T)\times\Omega,\\ u(T,x)&=h(x),\quad x\in\Omega,\end{split}$
where
$F_{\psi}(t,x,y,z)$ $:={\color[rgb]{0,0,0}\operatornamewithlimits{essinf}}_{a\in A}\Bigl{\{}\ell(t,x,a)+(f(t,x,a),z)-\lambda(t,x,a)y$
$\qquad\qquad+\int_{{\mathbb{R}}^{d}}\psi(t,x\otimes_{t}e)\lambda(t,x,a)\,Q(t,x,a,{\mathrm{d}}e)\Bigr{\}}.$
To define our notion of solution, we need the following path spaces. Given $(s_{0},x_{0})\in[0,T)\times\Omega$ and $z\in{\mathbb{R}}^{d}$ in addition to $\psi$, define
$\mathcal{X}(s_{0},x_{0})$ :=$:=\Big{\{}x\in\Omega:\,x=x_{0}\text{ on $[0,t_{0}]$, $x|_{[t_{0},T]}$ is absolutely continuous with}$
$\qquad\qquad\left|{x^{\prime}(t)}\right|\leq C_{f}(1+\sup_{s\leq t}\left|{x(s)}\right|)\text{ a.e.~{}on $(t_{0},T)$}\Big{\}}$
and
$\mathcal{Y}_{\psi}(s_{0},x_{0},y_{0},z)$ :=$:=\Big{\{}(x,y)\in\mathcal{X}(s_{0},x_{0})\times C([t_{0},T]):$
$\qquad\qquad y(t)=y_{0}+\int_{s_{0}}^{t}(x^{\prime}(s),z)-F_{\psi}(s,x,y(s),z)\,{\mathrm{d}}s\text{ on $[t_{0},T]$}\Big{\}}.$
Recall from section 3.1 that $[0,T]\times\Omega$ is considered to be equipped with the pseudo-metric defined in (3.1). Note that a function $u:[0,T]\times\Omega\to{\mathbb{R}}$ that is semicontinuous (with respect to that pseudo-metric) is non-anticipating, i.e., $u(t,x)=u(t,x_{\cdot\wedge t})$ for each $(t,x)\in[0,T]\times\Omega$.
@@
(i) $u$ is a minimax supersolution of ($\text{{TVP}}_{\psi}$) if $u\in\mathrm{LSC}([0,T]\times\Omega)$, if $u(T,\cdot)\geq h$ on $\Omega$, and if, for every $(s_{0},x_{0},z)\in[0,T)\times\Omega\times{\mathbb{R}}^{d}$, and every $y_{0}\geq u(s_{0},x_{0})$, there exists an $(x,y)\in\mathcal{Y}_{\psi}(s_{0},x_{0},y_{0},z)$ such that $y(t)\geq u(t,x)$ on $[s_{0},T]$.
(ii) $u$ is a minimax subsolution of ($\text{{TVP}}_{\psi}$) if $u\in\mathrm{USC}([0,T]\times\Omega)$, if $u(T,\cdot)\leq h$ on $\Omega$, and if, for every $(s_{0},x_{0},z)\in[0,T)\times\Omega\times{\mathbb{R}}^{d}$, and every $y_{0}\leq u(s_{0},x_{0})$, there exists an $(x,y)\in\mathcal{Y}_{\psi}(s_{0},x_{0},y_{0},z)$ such that $y(t)\leq u(t,x)$ on $[s_{0},T]$.
@@
(a) First, note that $G\psi_{0}$ is continuous (to see this, one can proceed as in the proof of Theorem 7.4 in [BK18JFA]). Thus, together with Theorem 96, 146-IV, in [DMprobPot], we can deduce that the maps
@@
Now, consider the terminal value problem
(TVP) $\begin{split}-\partial_{t}u(t,x)-F_{u}(t,x,u(t,x),\partial_{x}u(t,x))&=0,\quad(t,x)\in[0,T)\times\Omega,\\ u(T,x)&=h(x),\quad x\in\Omega.\end{split}$
Definition 5.10.
Let $u:[0,T]\times\Omega\to{\mathbb{R}}$.
(i) $u$ is a minimax supersolution of (TVP) if $u\in\mathrm{BLSA}([0,T]\times\Omega)$, $u=u(s,x)$ is Lipschitz in $x$, and $u$ is a minimax supersolution of ($\text{{TVP}}_{\psi}$) with $\psi=u$.
(ii) $u$ is a minimax subsolution of (TVP) if $u\in\mathrm{BLSA}([0,T]\times\Omega)$, $u=u(s,x)$ is Lipschitz in $x$, and $u$ is a minimax subsolution of ($\text{{TVP}}_{\psi}$) with $\psi=u$.
(iii) $u$ is a minimax solution of (TVP) if it is both a minimax supersolution and a minimax subsolution of (TVP).
Recall the fixed point $\bar{\psi}$ of the operator $G$ from Theorem 5.4.
@@
$\bar{\psi}$ is the unique minimax solution of (TVP).

### v2

@@ 5.2. Minimax solutions
(i) First, assume that $u$ is a classical solution of (5.1). We show that $u$ is a minimax supersolution of (5.1). To this end, fix $(s_{0},x_{0},z)\in[0,T)\times\Omega\times{\mathbb{R}}^{d}$. By Proposition 5.1 of [Lukoyanov03], there is a pair $(x,y)\in\mathcal{X}(s_{0},x_{0})\times C([s_{0},T])$ such that
(5.9) $\begin{split}y(t)&=u(s_{0},x_{0})+\int_{s_{0}}^{t}(x^{\prime}(s),z)-{\color[rgb]{0,0,0}H^{u}(s,x,u(s,x),z)}\,{\mathrm{d}}s\text{ and}\\ y(t)&=u(t,x)\end{split}$
for every $t\in[s_{0},T]$. We proceed by considering the following two cases.
Case 1. Let $y_{0}=u(s_{0},x_{0})$. By (5.9), we have $(x,y)\in\mathcal{Y}_{u}(s_{0},x_{0},y_{0},z)$ and $y(t)=u(t,x)$ on $s_{0},T]$.
Case 2. Let $y_{0}>u(s_{0},x_{0})$. Now, let $\tilde{y}$ be a solution of
(5.10) $\begin{split}\tilde{y}^{\prime}(t)&=x^{\prime}(t)\cdot z-{\color[rgb]{0,0,0}H^{u}(t,x,\tilde{y}(t),z)}\quad\text{a.e.~on $(s_{0},T)$},\\ \tilde{y}(s_{0})&=y_{0},\end{split}$
Then $(x,\tilde{y})\in\mathcal{Y}_{u}(s_{0},x_{0},y_{0},z)$. We claim that
(5.11) $\tilde{y}(t)>u(t,x)\text{ on $(s_{0},T)$.}$
For the sake of a contradiction, assume that (5.11) does not hold, i.e., there is a smallest time $s_{1}\in(s_{0},T)$ such that $\tilde{y}(s_{1})=u(t,s_{1})$ and $\tilde{y}(t)>u(t,x)$ on $(s_{0},s_{1})$. But then, by (5.9), (5.10), and (5.2), we have
$0$ $=\tilde{y}(s_{1})-u(s_{1},x)$
$=\tilde{y}(s_{0})-u(s_{0},x_{0})+\int_{s_{0}}^{s_{1}}\Bigl[{\color[rgb]{0,0,0}H^{u}(t,x,u(t,x),z)}{\color[rgb]{0,0,0}-H^{u}(t,x,\tilde{y}(t),z)}\Bigr]\,{\mathrm{d}}t$
$\geq\tilde{y}(s_{0})-u(s_{0},x_{0})$
$>0,$
which is a contradiction (cf. section 6.2 of [BK1]). Thus, (5.11) holds.
By Cases 1 and 2, we can conclude that $u$ is a minimax supersolution of (5.1). Similarly, one can show that $u$ is a minimax subsolution of (5.1). Hence, $u$ is a minimax solution of (5.1).
(ii) Now, assume that $u$ is a minimax solution of (5.1). Fix $(s_{0},x_{0})\in[0,T)\times\Omega$. Then there is an $(x,y)\in\mathcal{Y}_{u}(s_{0},x_{0},u(s_{0},x_{0}),\partial_{x}u(s_{0},x_{0}))$ such that $y(t)\geq u(t,x)$ on $s_{0},T]$, i.e., for every $t\in[s_{0},T]$, we have
$u(s_{0},x_{0})+\int_{s_{0}}^{t}\left[{\color[rgb]{0,0,0}x^{\prime}(s)\cdot\partial_{x}u(s_{0},x_{0})}-{\color[rgb]{0,0,0}H^{u}(s,x,y(s),\partial_{x}u(s_{0},x_{0}))}\right]\,{\mathrm{d}}s$
$\qquad\geq u(t,x),$
or, equivalently,
$\int_{s_{0}}^{t}\left[{\color[rgb]{0,0,0}x^{\prime}(s)\cdot\partial_{x}u(s_{0},x_{0})}-{\color[rgb]{0,0,0}H^{u}(s,x,y(s),\partial_{x}u(s_{0},x_{0}))}\right]\,{\mathrm{d}}s$
$\qquad\qquad\geq\int_{s_{0}}^{t}\left[\partial_{t}u(s,x)+x^{\prime}(s)\cdot\partial_{x}u(s,x)\right]\,{\mathrm{d}}s.$
Hence, by (5.8),
$-\partial_{t}u(s_{0},x_{0})-{\color[rgb]{0,0,0}H^{u}(s_{0},x_{0},u(s_{0},x_{0}),\partial_{x}u(s_{0},x_{0}))}\geq 0.$
Similarly, one can see that the opposite inequality holds as well, i.e., $u$ is a classical solution of (5.1). ∎
@@
(i) $u$ is a minimax supersolution of ($\text{{TVP}}_{\psi}$) if $u\in\mathrm{LSC}([0,T]\times\Omega)$, if $u(T,\cdot)\geq h$ on $\Omega$, and if, for every $(s_{0},x_{0},z)\in[0,T)\times\Omega\times{\mathbb{R}}^{d}$, and every $y_{0}\geq u(s_{0},x_{0})$, there exists an $(x,y)\in\mathcal{Y}_{\psi}(s_{0},x_{0},y_{0},z)$ such that $y(t)\geq u(t,x)$ for each $t\in[s_{0},T]$.
(ii) $u$ is a minimax subsolution of ($\text{{TVP}}_{\psi}$) if $u\in\mathrm{USC}([0,T]\times\Omega)$, if $u(T,\cdot)\leq h$ on $\Omega$, and if, for every $(s_{0},x_{0},z)\in[0,T)\times\Omega\times{\mathbb{R}}^{d}$, and every $y_{0}\leq u(s_{0},x_{0})$, there exists an $(x,y)\in\mathcal{Y}_{\psi}(s_{0},x_{0},y_{0},z)$ such that $y(t)\leq u(t,x)$ for each $t\in[s_{0},T]$.
(iii) $u$ is a minimax solution of ($\text{{TVP}}_{\psi}$) if $u$ is both a minimax supersolution and a minimax subsolution of ($\text{{TVP}}_{\psi}$).
The next result uses the operator $G$ from (5.6).
@@
(a) First, note that $G\psi_{0}$ is continuous (to see this, one can proceed as in the proof of Proposition 7.4 in [BK1]). Thus, together with Theorem 96, 146-IV, in [DMprobPot], we can deduce that the maps
@@
5.3. Properties of the operator $G$
Let us first recall the operator $G$ from (5.6): For each function $\psi\in\mathrm{BLSA}([0,T]\times\Omega)$ and $(s,x)\in[0,T]\times\Omega$,
(5.13) $\begin{split}&[G\psi](s,x)=\inf_{a\in\mathcal{A}}\Bigl\{\chi^{s,x,a}(T)\,h(\phi^{s,x,a})\\ &+\int_{s}^{T}\chi^{s,x,a}(t)\,\Bigl[\ell^{s,x,a}(t)+\int_{{\mathbb{R}}^{d}}\psi(t,\phi^{s,x,a}\otimes_{t}e)\,\lambda^{s,x,a}(t)\,Q^{s,x,a}(t,{\mathrm{d}}e)\Bigr]\,{\mathrm{d}}t\Bigr\}.\end{split}$
Remark 5.9.
Note that
$G:\mathrm{BLSA}([0,T]\times\Omega)\to\mathrm{BLSA}([0,T]\times\Omega).$
This follows from Theorem 96, 146-IV, in [DMprobPot], Lemma 4.5, Remark 4.6, Proposition 7.48 in [BS], and Proposition 7.47 in [BS].
Let us introduce several operators related to $G$. Given $s_{1}$, $s_{2}\in{\mathbb{R}}$ such that $0\leq s_{1}<s_{2}\leq T$, $\eta\in\mathrm{BLSA}([s_{2},T]\times\Omega)$, $a\in\mathcal{A}$, $\psi\in\mathrm{BLSA}([s_{1},s_{2}]\times\Omega)$, and $(s,x)\in[s_{1},s_{2}]\times\Omega$, define
(5.14) $\begin{split}&[(G_{s_{1},s_{2};\eta,a})\psi](s,x):=\chi^{s,x,a}(s_{2})\,\eta(s_{2},\phi^{s,x,a})+\int_{s}^{s_{2}}\Bigl[\chi^{s,x,a}(t)\,\ell^{s,x,a}(t)\\ &\qquad+\int_{{\mathbb{R}}^{d}}\psi(t,\phi^{s,x,a}\otimes_{t}e)\,\lambda^{s,x,a}(t)\chi^{s,x,a}(t)\,Q^{s,x,a}(t,{\mathrm{d}}e)\Bigr]\,{\mathrm{d}}t\qquad\text{and}\\ &[(G_{s_{1},s_{2};\eta})\psi](s,x):=\inf_{\tilde{a}\in\mathcal{A}}[(G_{s_{1},s_{2};\eta,\tilde{a}})\psi](s,x).\end{split}$
Note that $G_{s_{1},s_{2};\eta,a}$ and $G_{s_{1},s_{2};\eta}$ map bounded lower semi-analytic functions to bounded lower semi-analytic functions. This can be shown in the same way as the fact that $G$ maps bounded lower semi-analytic functions to bounded lower semi-analytic functions in Remark 5.9.
@@ 5.3. Properties of the operator $G$
The following holds:
(i) The operator $G$ has a unique fixed point $\bar{\psi}$ in $\mathrm{BLSA}([0,T]\times\Omega)$, which coincides with the value function $V$ from (4.6) on $(0,T]\times\Omega$.
@@
(i) Step 1 (construction of $\bar{\psi}$). Consider a partition $0=r_{0}<r_{1}<\ldots<r_{N}=T$ of $[0,T]$ with $r_{k}-r_{k-1}<\frac{1}{2}$ for all $k$. Then $G_{r_{N-1},r_{N};h}$ is a contraction on $\mathrm{BLSA}([r_{N_{1}},r_{N}]\times\Omega)$. To see this, note first that, for every $\alpha\in\mathcal{A}$, there is, by (5.14), Assumption 4.2, and the non-negativity of $\lambda$ (see Assumption 4.1), some probability measure $\mathbb{P}$}} with expected value $\mathbb{E}$$^{{\mathbb{P}}}$ such that, for each $\psi_{1}$, $\psi_{2}\in\mathrm{BLSA}([r_{N_{1}},r_{N}]\times\Omega)$,
$[G_{r_{N-1},r_{N};h,\alpha}\,\psi_{1}](s,x)-[G_{r_{N-1},r_{N};h,\alpha}\,\psi_{2}](s,x)$ $\leq\int_{r_{N-1}}^{r_{N}}{{\mathbb{E}}}^{{\mathbb{P}}}\lVert{\psi_{1}-\psi_{2}}\rVert_{\infty}\,{\mathrm{d}}t$
$\leq\frac{1}{2}\lVert{\psi_{1}-\psi_{2}}\rVert_{\infty}.$
Thus, for every $(s,x)\in[r_{N-1},r_{N}]\times\Omega$ and ${\varepsilon}>0$, there is an $\alpha^{\varepsilon}\in\mathcal{A}$ with
$[G_{r_{N-1},r_{N};h}\,\psi_{1}](s,x)-[G_{r_{N-1},r_{N};h}\,\psi_{2}](s,x)$
$\qquad\leq[G_{r_{N-1},r_{N};h}\,\psi_{1}](s,x)-[G_{r_{N-1},r_{N};h,\alpha^{\varepsilon}}\,\psi_{2}](s,x)+{\varepsilon}$
$\qquad\leq[G_{r_{N-1},r_{N};h,\alpha^{\varepsilon}}\,\psi_{1}](s,x)-[G_{r_{N-1},r_{N};h,\alpha^{\varepsilon}}\,\psi_{2}](s,x)+{\varepsilon}\leq\frac{1}{2}\lVert{\psi_{1}-\psi_{2}}\rVert_{\infty}+{\varepsilon}.$
Since $G_{r_{N-1},r_{N};h}$ is a contraction on $\mathrm{BLSA}([r_{N_{1}},r_{N}]\times\Omega)$, it has a unique fixed point $\bar{\psi}_{N}{\color[rgb]{0,0,0}\in\mathrm{BLSA}([r_{N_{1}},r_{N}]\times\Omega)}$.
Next, we show that the value function $V^{\prime}$ defined by (4.6) coincides with $\bar{\psi}_{N}$ on $r_{N-1},r_{N}]\times\Omega$. Formally, this would immediately follow from Theorem 4.17 and the uniqueness of the fixed point $\bar{\psi}$. Note however that at this point, we have not established yet that $V^{\prime}$ is lower semi-analytic. Thus we cannot apply $G$ (or the related operators defined by (5.14)) to $V^{\prime}$, as $G$ is an operator on $\mathrm{BLSA}([0,T]\times\Omega)$. Instead, we use $G^{\circ}$ from (4.18), which can be applied to $V^{\prime}$, and obtain
$\leq{\color[rgb]{0,0,0}\frac{1}{2}}\lVert{V^{\prime}|_{[r_{N-1},r_{N}]\times\Omega}-\bar{\psi}_{N}}\rVert_{\infty},$
i.e., $\bar{\psi}_{N}=V^{\prime}|_{[r_{N-1},r_{N}]\times\Omega}$ and thus $V^{\prime}$ is lower semi-analytic on $r_{N-1},r_{N}]\times\Omega$.
Similarly, $G_{r_{N-2},r_{N-1};\bar{\psi}_{N}}$ is a contraction on $\mathrm{BLSA}([r_{N-2},r_{N-1}]\times\Omega)$ and thus has a unique fixed point $\bar{\psi}_{N-1}{\color[rgb]{0,0,0}\in\mathrm{BLSA}([r_{N-2},r_{N-1}]\times\Omega)}$, which coincides with $V^{\prime}$ on $r_{N-2},r_{N-1}]\times\Omega$ thanks to Theorem 4.17.
Iterating this process yields a finite sequence $(\bar{\psi})_{k=1}^{N}$ with the properties that each $\bar{\psi}_{k}$ is the unique fixed point of $G_{r_{k-1},r_{k};\bar{\psi}_{k+1}}$ on $\mathrm{BLSA}([r_{k-1},r_{k}]\times\Omega)$, where we set $\bar{\psi}_{N+1}(s,x):=h(x)$, and that $\bar{\psi}_{k}$ coincides with $V^{\prime}$ on $r_{k-1},r_{k}]\times\Omega$ if $k>2$ and on $(0,r_{1}]\times\Omega$ if $k=1$.
Next, define $\bar{\psi}:[0,T]\times\Omega\to{\mathbb{R}}$ by
Step 2 ($\bar{\psi}$ is a fixed point of $G$). Consider the function $\bar{v}:[0,T]\times\Omega\to{\mathbb{R}}$ defined by $\bar{v}(s,x):=[G\bar{\psi}](s,x)$. Then $\bar{v}=\bar{\psi}_{N}$ on $r_{N-1},r_{N}]\times\Omega$.
$\bar{v}(s,x){\color[rgb]{0,0,0}=[(G_{0,T;h})\bar{\psi}](s,x)}=[(G_{r_{N-2},r_{N-1};\bar{v}})\bar{\psi}_{N-1}](s,x)=\bar{\psi}_{N-1}(s,x).$
Iterating this procedure yields $\bar{v}=\bar{\psi}$, i.e., $\bar{\psi}$ is a fixed point of $G$.
Another way to verify the desired fixed point property of $\bar{\psi}$ would be to just note that, by Step 1, $\bar{\psi}$ coincides with $V$ and is lower semi-analytic. Then, together with Corollary 4.20, $V=GV$ follows. However, note this argument relies on the rather technical section 4.4 unlike the argument in the preceding paragraphs.
Step 3 ($\bar{\psi}$ is the unique fixed point of $G$). Let $\bar{\psi}^{\prime}$ be a fixed point of $G$. On $r_{N-1},r_{N}]\times\Omega$, we have $\bar{\psi}^{\prime}=G\bar{\psi}^{\prime}=(G_{r_{N-1},r_{N};h})\bar{\psi}^{\prime}$. Thus $\bar{\psi}^{\prime}$ coincides on $r_{N-1},r_{N}]\times\Omega$ with $\bar{\psi}_{N}$, the unique fixed point of $G_{r_{N-1},r_{N};h}$.
Hence, $\bar{\psi}^{\prime}$ coincides on $r_{N-2},r_{N-1}]\times\Omega$ with $\bar{\psi}_{N-1}$, the unique fixed point of $G_{r_{N-2},r_{N-1};\bar{\psi}_{N}}$. Iterating this process yields $\bar{\psi}^{\prime}=\bar{\psi}$.
Next, we show that the fixed point $\bar{\psi}$ is sufficiently regular. To this end, we need the following auxiliary result, whose proof is relegated to subsection B.2.
Lemma 5.11.
for all $s\in[s_{1},s_{2}]$, $x$, $\tilde{x}\in\Omega$. Let $\lVert{\eta}\rVert_{\infty}\leq\lVert{\bar{\psi}}\rVert_{\infty}$. Recall $G_{s_{1},s_{2};\eta}$ from (5.14). Then
for all $(s,x,\tilde{x})\in[s_{1},s_{2}]\times\Omega\times\Omega$, where
Theorem 5.12.
There exists a constant $L_{\bar{\psi}}\in{\mathbb{R}}_{+}$ such that, for every $s\in[0,T]$ and every $x$, $\tilde{x}\in\Omega$, we have
Proof.
By mathematical induction,
(5.18) $c_{n}(c^{\prime})\leq 2c^{\prime}{\mathrm{e}}^{L_{f}T}+1\text{ for all $n\in{\mathbb{N}}$.}$
$r_{j-1}=s_{j-1,0}<s_{j-1,1}<\ldots<s_{j-1,M}=r_{j}$
and, for each $k\in\{0,\ldots,M-1\}$,
(5.19) $s_{j,k+1}-s_{j,k}\leq\frac{1}{12\check{L}{\color[rgb]{0,0,0}\left(1+\sup_{n\in{\mathbb{N}}_{0}}\lVert{(G^{n}_{r_{j-1},r_{j};\bar{\psi}})\psi_{0}}\rVert_{\infty}\right)}}=:\check{m}_{j}.$
Step 1. Let $n\in{\mathbb{N}}_{0}$. By (5.19) and Lemma 5.11 applied to $(G^{n}_{r_{N-1},r_{N};\bar{\psi}})\psi_{0}]$ restricted to $s_{N-1,M-1},s_{N-1,M}]\times\Omega$, $c_{n}(L_{f})$, $L_{f}$, $h$, $s_{N-1,M-1}$, $s_{N-1,M}$ in place of $\psi$, $c$, $c^{\prime}$, $\eta$, $s_{1}$, $s_{2}$ and noting that
we can see that
[…]
for all $s\in[s_{N-1,M-1},s_{N-1,M}]$, $x$, $\tilde{x}\in\Omega$.
By Theorem 5.10 (ii), letting $n\to\infty$ yields (5.17) on $s_{N-1,M-1},s_{N-1,M}]\times\Omega\times\Omega$ with $c^{\prime}_{N}:=2L_{f}{\mathrm{e}}^{L_{f}T}+1$ in place of $L_{\bar{\psi}}$.
Step 2. Let $n\in{\mathbb{N}}_{0}$. Without loss of generality, assume that $r_{N-1}\leq s_{N-1,M-2}$. Consequently, by (5.19) and Lemma 5.11 applied to $(G^{n}_{r_{N-1},r_{N};\bar{\psi}})\psi]$ restricted to $s_{N-1,M-2},s_{N-1,M-1}]\times\Omega$, $c_{n}(c^{\prime}_{N})$, $c^{\prime}_{N}$, $\bar{\psi}$, $s_{N-1,M-2}$, $s_{N-1,M-1}$ in place of $\psi$, $c$, $c^{\prime}$, $\eta$, $s_{1}$, $s_{2}$, we can with Step 1 of this proof deduce that
[…]
for all $s\in[s_{N-1,M-2},s_{N-1,M-1}]$, $x$, $\tilde{x}\in\Omega$. By Theorem 5.10 (ii), letting $n\to\infty$ yields (5.17) on $s_{N-1,M-2},s_{N-1,M-1}]\times\Omega\times\Omega$ with $2c^{\prime}_{N}{\mathrm{e}}^{L_{f}T}+1$ in place of $L_{\bar{\psi}}$.
Step 3. Repeating the procedure of Step 2 finitely many times and making the obvious adjustments every time we pass to a different interval $r_{k-1},r_{k}]$, we obtain the existence of some constant $L_{\bar{\psi}}\in{\mathbb{R}}_{+}$ such that (5.17) holds on $0,T]\times\Omega\times\Omega$. ∎
5.4. Existence and uniqueness
Recall from Theorem 5.10 the fixed point $\bar{\psi}$ of the operator $G$.
Theorem 5.13.
$\bar{\psi}$ is the unique minimax solution of (5.1).
Proof.